40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Wed Oct 30 01:56:38 2013, tara, DailyProgress, optic, table work oldPBS.jpgnewPBS.jpg
    Reply  Thu Oct 31 00:02:17 2013, tara, DailyProgress, Electronics Equipment, TTFSS offset.pngD040105C.png
       Reply  Thu Oct 31 18:10:07 2013, Evan, DailyProgress, Electronics Equipment, TTFSS 
       Reply  Sun Nov 3 19:37:32 2013, rana, DailyProgress, Electronics Equipment, TTFSS 
          Reply  Tue Nov 5 01:11:15 2013, tara, DailyProgress, Electronics Equipment, TTFSS 
             Reply  Wed Nov 6 01:14:58 2013, tara, DailyProgress, Electronics Equipment, TTFSS beat_2013_11_05.pngbeat_2013_11_05.fig
                Reply  Thu Nov 7 05:08:13 2013, tara, DailyProgress, NoiseBudget, photothermal noise in SiO2/Ta2O5 beat_2013_11_07.pngbeat_2013_11_07.fig
                   Reply  Wed Nov 20 18:19:01 2013, tara, DailyProgress, NoiseBudget, photothermal noise in SiO2/Ta2O5 beat_compare.pngbeat_compare.fig
Message ID: 1378     Entry time: Thu Oct 31 18:10:07 2013     In reply to: 1377
Author: Evan 
Type: DailyProgress 
Category: Electronics Equipment 
Subject: TTFSS 

Tara and I spent some time looking at the TTFSS boards. The offset issues appear to be caused by bad choice of offset knob on the TTFSS interface boards. Previously, Tara and I had used the offset knob to null RFAM-induced offsets in the north PDH signal. The current thinking is that when Tara re-optimized the electro-optic elements on the north path, the RFAM-induced offset changed and was therefore no longer nulled.

We have now returned the offset knobs to their optimal values as follows: we set each TTFSS to use the TEST SMA input rather than the LO/PD + mixer input, we applied a 50 Ω terminator to this input, and we then watched TP4 (after the common VGA) on a scope while adjusting the offset knobs. The optimal knob positions are 526 for the south TTFSS and 506 for the north TTFSS. Varying the common gain causes the DC offset on TP4 to change only slightly (it stays within ±5 mV for both north and south). Varying the fast gain causes the DC offset on TP17 (after the fast VGA) to vary as well; on south, this also appears to stay within ±5 mV of zero, but on north it is as high as 20 mV when the common and fast gains are turned all the way up. However, since these VGAs are each +30 dB at maximum gain, this means that the offset referred to TTFSS OUT1 is more like 20 µV, which is negligible compared to an error signal that is something like 1 Vpp.

ELOG V3.1.3-